OpenStudy (anonymous):

\[q^2-64/q^2-16q+64\] simplify by removing factors of 1.

OpenStudy (anonymous):

\[(q^2 + 8)/(q^2 - 8)\]

OpenStudy (anonymous):

it says wrong. that you for trying though!

OpenStudy (anonymous):

MagRat almost right. but its (q+8)/(q-8)

OpenStudy (anonymous):

\[\frac{q^2-64}{q^2-16+64}\] \[\frac{(q+8)(q-8)}{(q-8)(q-8)}\] \[\frac{q+8}{q-8}\] is the method

OpenStudy (anonymous):

thank you!